OpenStudy (anonymous):

I know the answer is 4 I just don't know how they got it.

OpenStudy (anonymous):

\[\sqrt{16-2b}=\sqrt{2b}\]

OpenStudy (anonymous):

Wouldn't the -2b end up canceling the positive 2b and then you'd just be left with the square root of 16 which is 4?

OpenStudy (anonymous):

No, remember that signs change when switching sides of the equation. \[\sqrt{16-2b} = \sqrt{2b}\] \[\sqrt{16} - \sqrt{2b}\ = \sqrt{2b}\] \[\sqrt{16} = \sqrt{2b}\ +\sqrt{2b}\] \[\sqrt{16} = \sqrt{2b+2b}\] \[\sqrt{16} = \sqrt{4b}\] or b=4

OpenStudy (anonymous):

Thanks!(: